Brenda and Trent are friends. Brenda is 24. The LCM of their age is 120 and the GCF of their age is 8. How old is Trent.
Answer by mathmaven53(29)

Factor Brenda's age into 24 = 3*8
Factor 120 into 120 = 8*3*5
Trent can't be 120 since GCF of their ages would be 3*8
Trent can't be 8*3 since GCF of their ages would be 8*3
Let Trent be 8*5 = 40
GCF(24,40) = 8
24 = 3*8
40 = 5*8
5*24 = 3*5*8=120
3*40 = 3*5*8 =120
LCM(24,40)=120
So Trent is 40